# PEL-NAS: Search Space Partitioned Architecture Prompt Co-Evolutionary LLM-driven Hardware-Aware Neural Architecture Search

### v1

@@ 1 Introduction
Figure 1: Comparison of three generation strategies on HW-NAS-Bench (Edge GPU, CIFAR-10): normal prompt (orange), latency-optimized prompt (blue), and PEL-NAS (green). Latency-optimized prompting increases coverage compared to standard prompting but still leaves gaps, while PEL-NAS achieves near-complete coverage across latency ranges.
Recently, Large Language Models (LLMs) offer a promising training-free alternative for discovering neural architectures Achiam et al. (2023). However, applying an LLM directly to the vast HW-NAS search space raises two challenges. First, we observe the exploration bias issue, which is analogous to the mode collapse issue in generative models Shumailov et al. (2024); Kossale et al. (2022); Zhang et al. (2025). Specifically, the LLM tends to repeatedly generate safe and familiar architectural patterns within limited search space, without fully exploring the full search space. Figure 1 compares three generation strategies on HW-NAS-Bench (Edge GPU, CIFAR-10). In (a) Normal prompt, we give only a plain task description including target device and dataset and ask the LLM to propose an architecture. The LLM then concentrates in a small area with limited coverage of the latency range. In (b) Latency-optimized prompt, we add an explicit hint to aim for diverse latencies and pass back the previous round’s accuracy and latency to the LLM. The results shift toward lower latency but the coverage remains uneven. The number of low-latency architectures attempted by LLM is still small and not competitive. This motivates the development of a strategy that can further expand search space. Second, most existing LLM approaches rely on static prompts, lacking a mechanism to accumulate knowledge from past evaluations. Without this feedback loop, the LLM cannot refine its design rules over generations, which slows progress toward the true Pareto front.
@@ 3 Methodology
Figure 2: The PEL-NAS framework. The search space is partitioned into complexity-based niches, where an LLM performs parallel evolutionary search. The individual results are then aggregated to form the final, complete Pareto front, mitigating exploration bias.
@@ 3.1 Complexity-Driven Search Space Partitioning
Figure 3: Analysis of the HW-NAS-Bench search space. The distribution of total parameters exhibits clear clustering, where each cluster corresponds to a specific number of nor_conv_3x3.
@@ 3.2 LLM-Powered Partitioned Co-Evolution of Prompts and Architectures
Figure 4: The Co-evolve Prompt Generator in PEL-NAS. The LLM first acts as a reasoning engine, updating a Knowledge Base by analyzing prior results. This learned knowledge then informs the LLM’s second role as an expert architect, where it generates new, rationale-driven architectures under specific constraints, creating a self-optimizing search process.
@@ 4.1 Experimental Setup
  • Hypervolume (HV): This metric measures the overall coverage of the discovered front. It rewards fronts that contain a wide variety of solutions that are both highly accurate and fast. Formally, given a reference point $r$ that is dominated by all solutions in the front $S$, the HV is the volume of the region bounded by the front and the reference point:
@@
    A larger HV is better, indicating a more complete and higher-quality front.
  • Inverted Generational Distance (IGD): This metric measures the closeness or fidelity of our discovered front to the true, perfect front. It essentially answers the question: On average, how far away is each theoretically perfect solution from the nearest solution we actually found? It is defined as:
@@
    where $d(\cdot,\cdot)$ is the Euclidean distance. A lower IGD is better, signifying a more accurate approximation of the true optimal front.
@@ 4.2 Main Results
Table 2: Comparison of selected top structures of HW-NAS-Bench on CIFAR-10. Acc.=Top-1 Accuracy, Lat.=Latency
@@
Analysis of Discovered Architectures of HW-NAS-Bench on CIFAR-10. Beyond the overall front quality, the individual architectures in Table 2 highlight the practical value of our method. PEL-NAS not only finds models with state-of-the-art accuracy, matching the performance of costly supernet-based methods, but also excels in the low-latency domain where other approaches falter. Crucially, it discovers the undisputed fastest architecture for each hardware target. For example, it identifies a model with a latency of just 1.78ms on the Edge GPU and 1.65ms on the FPGA—outperforming the fastest competitor, PRP-NAS-BL, by over 22% and 54% respectively. This proves its superior ability to explore the full spectrum of trade-offs and deliver a truly comprehensive set of optimal solutions.
Pareto Front Quality Evaluation with HV and IGD. The core quantitative results in Table 3 compare the discovered Pareto fronts using HV and IGD. Across all three datasets and six hardware targets, PEL-NAS consistently and significantly outperforms all baselines. PEL-NAS achieves higher HV scores and the lower IGD scores compared with baselines. For example, on CIFAR-10, PEL-NAS can achieve up to 80.6% higher HV and 53.6% lower IGD compared with non-constrained LLM Method. On CIFAR-100, PEL-NAS outperforms LLMatic, FairNAS, PRP-NAS, by 46.5%, 5.7%, 17.4% in HV, and by 34.9%, 6.1%, 18.6% in IGD, in average respectively. These observations further confirm that the front discovered by PEL-NAS is not only larger in volume but also much closer to the true optimal front. The experimental results also demonstrate that our complexity-driven partitioning strategy is highly effective in mitigating the LLM’s generative bias and enabling a more complete and diverse exploration of the search space.
@@ 4.4 Generalizability on Vision Transformer Search Spaces
To create a realistic hardware-aware scenario, we profile the latency of each candidate architecture directly on our target device, a single NVIDIA A6000 GPU. We then apply the core principle of PEL-NAS—complexity-driven partitioning. Our analysis of the ViT architecture (see Appendix D for a detailed breakdown) reveals that computational complexity, a strong proxy for latency, is dominated by two key parameters: Embed Dim (quadratic impact, $O(D^{2})$) and Depth Num (linear impact, $O(L)$). These parameters govern the scale of the MLP and the number of blocks, respectively, making them the most influential factors. We therefore partition the search space into niches based on discrete ranges of Embedding Dimension and Depth Number, enabling the LLM to efficiently explore trade-offs within structurally similar architectural families. The results, depicted in Figure 5 and detailed in Table 6, underscore the efficacy of our approach. PEL-NAS successfully identifies a dominant Pareto front, discovering architectures with superior accuracy-latency trade-offs.
@@ 5 Conclusion
Appendix A LLM Usage Disclosure
We used large language models (LLMs) in two ways. (1) Method component: within PEL-NAS, an LLM serves as a co-evolutionary operator (Section 3.2) to generate candidates with rationale under niche constraints. (2) Writing assistance: we additionally used LLMs for minor editing (grammar, wording, and clarity). No generated text was used as scientific evidence without verification, and all experiments are fully reproducible from the described algorithms and released code.
Appendix B Algorithm
@@
Appendix C Result of Ablation Study on all Datasets and Devices
@@
Figure 6: Results of the ablation study on CIFAR-10 across six hardware devices. Each subplot compares the Pareto fronts discovered by our full model (PEL-NAS) against its three ablated versions. The consistent dominance of the full PEL-NAS model demonstrates that each component is crucial for discovering the optimal trade-off between accuracy and latency.
Figure 7: Results of the ablation study on CIFAR-100 across six hardware devices.
Figure 8: Results of the ablation study on ImageNet16-120 across six hardware devices.
Appendix D Computational Complexity Analysis of the Vision Transformer Search Space
@@
Appendix E Analysis of LLM Exploration Bias
@@
Figure 9: LLM’s mode collapse in NAS persists despite prompt engineering. The figure shows the Pareto fronts discovered by an unpartitioned LLM-driven method, providing clear visual evidence of mode collapse. The LLM-generated architectures are highly clustered in a narrow region of the performance-latency space, resulting in a sparse and incomplete Pareto front that finds far fewer non-dominated solutions. This failure to explore—i.e., mode collapse—occurs even when the LLM is explicitly prompted to target diverse latencies, powerfully demonstrating the need for a more structural intervention, like our proposed partitioning strategy, to effectively guide the generative process.
Appendix F LLM Prompt Templates and Co-evolution Process

### v2

@@ 1 Introduction
Figure 1: Comparison of three generation strategies on HW-NAS-Bench (Edge GPU, CIFAR-10): normal prompt (orange), latency-optimized prompt (blue), and PEL-NAS (green). Latency-optimized prompting increases coverage compared to standard prompting but still leaves gaps, while PEL-NAS achieves near-complete coverage across latency ranges
Recently, Large Language Models (LLMs) offer a promising training-free alternative for discovering neural architectures Achiam et al. (2023). However, applying an LLM directly to the vast HW-NAS search space raises two challenges. First, we observe the exploration bias issue, which is analogous to the mode collapse issue in generative models Shumailov et al. (2024); Kossale et al. (2022); Zhang et al. (2025). Specifically, the LLM tends to repeatedly generate safe and familiar architectural patterns within limited search space, without fully exploring the full search space. Figure 1 compares three generation strategies on HW-NAS-Bench (Edge GPU, CIFAR-10). In (a) Normal prompt, we give only a plain task description including target device and dataset and ask the LLM to propose an architecture. The LLM then concentrates in a small search space with limited coverage of the latency range. In (b) Latency-optimized prompt, we add an explicit hint to aim for diverse latencies and pass back the accuracy and latency of the previous round to the LLM. The results shift toward lower latency, but the coverage remains uneven. The number of low-latency architectures attempted by LLM is still small and not competitive. This motivates the development of a strategy that can further expand search space. Second, most existing LLM approaches Zheng et al. (2023); Nasir et al. (2024) rely on static prompts , lacking a mechanism to accumulate knowledge from past evaluations. Without this feedback loop, the LLM cannot refine its design rules over generations, which slows progress toward the true Pareto front.
@@ 3 Methodology
Figure 2: The PEL-NAS framework. The search space is partitioned into complexity-based niches, where an LLM performs parallel evolutionary search. The individual results are then aggregated to form the final, complete Pareto front, mitigating exploration bias
@@ 3.1 Complexity-Driven Search Space Partitioning
Figure 3: Analysis of the HW-NAS-Bench search space. The distribution of total parameters exhibits clear clustering, where each cluster corresponds to a specific number of nor_conv_3x3
@@ 3.2 LLM-Powered Partitioned Co-Evolution of Prompts and Architectures
Figure 4: The Co-evolve Prompt Generator in PEL-NAS. The LLM first acts as a reasoning engine, updating a Knowledge Base by analyzing prior results. This learned knowledge then informs the LLM’s second role as an expert architect, where it generates new, rationale-driven architectures under specific constraints, creating a self-optimizing search process
@@ 4.1 Experimental Setup
  • Hypervolume (HV): This metric evaluates the diversity and quality of our approach by measuring the volume of the multi-objective space dominated by the discovered front $S$ relative to a reference point $r=(r_{1},r_{2},\dots,r_{m})$ dominated by all solutions in $S$, where $S$ is a set of solution points, $m$ is the number of objectives. A larger volume indicates a front with a set of solutions that are superior and more diverse. Formally, it is defined as:
@@
    where $s=(s_{1},\dots,s_{m})$ is a solution point in front $S$ and $s_{i}$ is the value of the $i$-th objective. In our experimental setting, m is selected as 2 for accuracy and latency.
  • Inverted Generational Distance (IGD): This metric primarily measures fidelity by calculating the average Euclidean distance from each solution in the true Pareto front ($P^{*}$) to its nearest neighbor in our discovered front ($S$). It directly quantifies how faithfully the discovered front approximates the ground-truth optimal set. It is defined as:
@@
    where $|P^{*}|$ is the number of solutions in the true front. A lower IGD score is better, indicating a closer approximation to the true optimum.
@@ 4.2 Main Results
Table 2: Comparison of selected top structures of HW-NAS-Bench on CIFAR-10. Acc. and Lat. refer to Top-1 Accuracy and Latency, respectively
@@
Analysis of Discovered Architectures of HW-NAS-Bench on CIFAR-10. Beyond the overall front quality, the individual architectures in Table 2 highlight the practical value of our method. PEL-NAS not only finds models with state-of-the-art accuracy, matching the performance of costly supernet-based methods, but also excels in the low-latency domain where other approaches falter. Crucially, it discovers the undisputed fastest architecture for each hardware target. For example, it identifies a model with a latency of just 1.78ms on the Edge GPU and 1.65ms on the FPGA which outperforming PRP-NAS-BL by over 22% and 54% respectively. This proves its superior ability to explore the full spectrum of trade-offs and deliver a truly comprehensive set of optimal solutions.
Pareto Front Quality Evaluation with HV and IGD. The core quantitative results in Table 3 compare the discovered Pareto fronts using HV and IGD. Across all three datasets and six hardware targets, PEL-NAS consistently and significantly outperforms all baselines. PEL-NAS achieves higher HV scores and lower IGD scores compared to baselines. For example, on CIFAR-10, PEL-NAS can achieve up to 80.6% higher HV and 53.6% lower IGD compared to the non-constrained LLM method. On CIFAR-100, PEL-NAS outperforms LLMatic, FairNAS, PRP-NAS, by 46.5%, 5.7%, 17.4% in HV, and by 34.9%, 6.1%, 18.6% in IGD, in average respectively. These observations further confirm that the front discovered by PEL-NAS is not only larger in volume but also much closer to the true optimal front. The experimental results also demonstrate that our complexity-driven partitioning strategy is highly effective in mitigating the LLM’s generative bias and enabling a more complete and diverse exploration of the search space.
@@ 4.4 Generalizability on Vision Transformer Search Spaces
To create a realistic hardware-aware scenario, we profile the latency of each candidate architecture directly on our target device, a single NVIDIA A6000 GPU. We then apply the core principle of PEL-NAS—complexity-driven partitioning. Our analysis of the ViT architecture (see Appendix C for a detailed breakdown) reveals that computational complexity, a strong proxy for latency, is dominated by two key parameters: Embed Dim (quadratic impact, $O(D^{2})$) and Depth Num (linear impact, $O(L)$). These parameters govern the scale of the MLP and the number of blocks, respectively, making them the most influential factors. We therefore partition the search space into niches based on discrete ranges of Embedding Dimension and Depth Number, enabling the LLM to efficiently explore trade-offs within structurally similar architectural families. The results, depicted in Figure 5 and detailed in Table 6, underscore the efficacy of our approach. PEL-NAS successfully identifies a dominant Pareto front, discovering architectures with superior accuracy-latency trade-offs.
@@ 5 Conclusion
Appendix A Algorithm
@@
Appendix B Result of Ablation Study on all Datasets and Devices
@@
Figure 6: Results of the ablation study on CIFAR-10 across six hardware devices. Each subplot compares the Pareto fronts discovered by our full model (PEL-NAS) against its three ablated versions. The consistent dominance of the full PEL-NAS model demonstrates that each component is crucial for discovering the optimal trade-off between accuracy and latency
Figure 7: Results of the ablation study on CIFAR-100 across six hardware devices
Figure 8: Results of the ablation study on ImageNet16-120 across six hardware devices
Appendix C Computational Complexity Analysis of the Vision Transformer Search Space
@@
Appendix D Analysis of LLM Exploration Bias
@@
Figure 9: LLM’s mode collapse in NAS persists despite prompt engineering. The figure shows the Pareto fronts discovered by an unpartitioned LLM-driven method, providing clear visual evidence of mode collapse. The LLM-generated architectures are highly clustered in a narrow region of the performance-latency space, resulting in a sparse and incomplete Pareto front that finds far fewer non-dominated solutions. This failure to explore—i.e., mode collapse—occurs even when the LLM is explicitly prompted to target diverse latencies, powerfully demonstrating the need for a more structural intervention, like our proposed partitioning strategy, to effectively guide the generative process
Appendix E LLM Prompt Templates and Co-evolution Process